OpenStudy (anonymous):

Factorise fully (n^2 - a^2) - (n - a)^2

OpenStudy (anonymous):

u can write (n^2 - a^2) =(n-a)(n+a) next take (n-a) as common n proceed

OpenStudy (lgbasallote):

\[n^2 - a^2 - (n-a)^2 \implies n^2- a^2 - (n^2 -2an + a^2) \implies n^2 - a^2 - n^2 + 2an - a^2\] does that help?

OpenStudy (anonymous):

use formula (a2-b2)=(a+b)(a-b )

OpenStudy (anonymous):

sai1234 - thanks that was simple! lgbasallote - sorry it was a bit confusing. thanks anyway!